### v1

# Building an Open AIBOM Standard in the Wild

Abstract.
Modern software engineering increasingly relies on open, community-driven standards, yet how such standards are created in fast-evolving domains like AI-powered systems remains underexplored. This paper presents a detailed experience report on the development of the AI Bill of Materials (AIBOM) specification, an extension of the ISO/IEC 5962:2021 Software Package Data Exchange (SPDX) software bill of materials (SBOM) standard, which captures AI components such as datasets and iterative training artifacts. Framed through the lens of Action Research (AR), we document a global, multi-stakeholder effort involving over 90 contributors and structured AR cycles. The resulting specification was validated through four complementary approaches: alignment with major regulations and ethical standards (e.g., EU AI Act and IEEE 7000 standards), systematic mapping to six industry use cases, semi-structured practitioner interviews, and an industrial case study. Beyond delivering a validated artefact, our paper documents the process of building the AIBOM specification “in the wild,” and reflects on how it aligns with the AR cycle, and distills lessons that can inform future standardization efforts in the software engineering community.
standardization, AIBOM, SBOM, software bill of materials, Action Research, traceability, transparency, AI governance, AI compliance, trustworthy AI, software supply chain security, SPDX, Software Package Data Exchange, open source community
conference: IEEE/ACM 48th International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP); April 12–18, 2026; Rio De Janeiro, Brazilcopyright: acmcopyrightbooktitle: IEEE/ACM 48th International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP), April 12–18, 2026, Rio De Janeiro, Brazilccs: Computing methodologies Machine learningccs: Software and its engineering Software development techniques
1. Introduction
Standards play a critical role in Software Engineering (SE) by ensuring consistency, interoperability, reliability, and trustworthiness across systems and end users (Fenton and Neil, 2002). They underpin much of the discipline. For instance, ISO/IEC 5962:2021, Software Package Data eXchange (SPDX), offers a shared vocabulary for documenting software supply chains (SSCs) that has been widely adopted to assess vulnerabilities and compliance in received software.
Regulators also rely on standards as mechanisms for compliance. In the EU, conformity with the Medical Device Regulation can be shown through harmonised standards such as ISO 13485 (quality management) and IEC 62304 (software lifecycle processes). In the U.S., Software Bill of Material (SBOM) standards like SPDX and CycloneDX are explicitly referenced in federal cybersecurity guidance (of Commerce, 2021; Cybersecurity and Infrastructure Security Agency, 2025) and medical device regulation. These SBOM standards have spurred a plethora of recent work that has examined SBOM adoption, proposed taxonomies, and developed extensions for vulnerability management (Stalnaker et al., 2024; Xia et al., 2024), which underscores their growing importance.
These SBOM standards target traditional software and cannot meaningfully capture AI-specific artefacts such as datasets, trained models, fine-tuned checkpoints, and data pipelines that dominate modern systems (Xia et al., 2023a). To address this limitation, the AI community has introduced transparency frameworks such as model cards (Mitchell et al., 2019), datasheets (Gebru et al., 2021), and factsheets (et al., 2019), as well as machine-readable formats like Croissant (et al., 2024). Yet, these approaches often isolate AI-specific details from the broader software engineering context in which practitioners build and deploy systems. Sculley et al. (Sculley et al., 2015) showed in their seminal work on hidden technical debt that AI components typically form only a small fraction of much larger systems, underscoring the need to document them in situ within the broader software context. Practitioners thus face a critical gap: they lack a unified, standard mechanism to describe AI components as first-class citizens within the software supply chain (Rajbahadur et al., 2021; Hassan et al., 2024).
Our paper presents an experience report on tackling this challenge through the creation of the AI Bill of Materials (AIBOM) specification, an extension to the SPDX SBOM standard. Our SPDX 3.0 extension comprises 36 new fields that treat datasets, models, and their provenance as first-class supply-chain elements to address the trustworthiness challenges of AI systems. Our work makes a distinct contribution by showing how AR can be adapted from its traditional use within single organization (Baskerville, 1999) to guide a global, multi-stakeholder standardization effort “in the wild.” Through this lens, our AIBOM specification (which extends the SBOM standard to represent AI systems) itself emerges as a validated artifact of a disciplined, iterative process. Over 90 contributors participated in diagnose-design-evaluate-reflect cycles, providing a replicable blueprint for developing standards in fast-moving domains in SE, while delivering a practical, community-driven specification.
AIBOM specification has since landed in the SPDX standard, been piloted in industrial settings, and informed regulatory discussions. While the technical specification and schema live in our whitepaper (Bennet et al., 2024), this paper reports our experience in open standardization, detailing our governance, cadence, and pitfalls. We also demonstrate AIBOM specification’s practical utility using four complementary validation methods: (1) regulatory/standards alignment (EU AI Act, US and EU medical-device guidance, IEEE 7000 series) to support compliance practice; (2) systematic mapping to industry use cases (compliance, risk management, supply-chain governance) with coverage analysis; (3) practitioner interviews (n=10) across roles confirming field relevance and surfacing gaps; and (4) a multinational industrial case study showing AIBOM covers ethics/legal checklists, generates portions of model cards, and enables partial automation for third-party AI assets. Together, these results position AIBOM specification as part of an actionable standard and our process as a blueprint for creating standards in fast-moving, multi-stakeholder SE domains.
Contributions. We report how we created AIBOM specification through an open, global AR process and why this approach worked in practice. Our aim is not to claim that SPDX AIBOM specification is the definitive AIBOM solution as credible alternatives such as CycloneDX exist and continue to evolve. Specifically:
  • AR at scale. We demonstrate how AR, traditionally used within single organizations (Baskerville, 1999), can be adapted to guide a global, multi-stakeholder standardization effort “in the wild.” Our account documents governance design, release cadence, decision traceability, and iterative cycles that extended SPDX SBOM standard with 36 AI and Dataset specific fields.
  • Validated artifact. We present AIBOM and demonstrate its practicality via four complementary validations.
  • Lessons and blueprint. We distill actionable patterns and lessons learned for creating standards in fast-moving, multi-stakeholder domains.
2. Background
This section contextualises our work by examining the formal processes for creating software engineering standards, the technical foundation of the SBOM that requires extension.
2.1. Development and Extension of Software Standards
Standards in SE. Standards are essential in SE as they provide common processes and terminology that minimize ambiguity, enhance interoperability, and ensure the reliable delivery of high-quality systems (Tsui et al., 2022). Broadly, the current standards can be grouped into key categories ranging from lifecycle processes and quality assessment to testing, supply-chain security, and emerging AI-specific considerations, as summarized in Table 1. These standards underpin the structured development, deployment, and governance of modern enterprise software, forming the foundation for consistent engineering practices across diverse domains and organizations.
Table 1. Representative categories of software engineering standards and their scope.
Category Examples Purpose / Scope
Lifecycle process ISO/IEC/IEEE 12207, 15288 Define processes for development, maintenance, and complete system lifecycles.
Quality ISO/IEC 25010 Provide models for evaluating core software quality attributes.
Testing & verification ISO/IEC/IEEE 29119 Standardise testing methods, documentation, and reporting practices.
Supply chain & compliance ISO 28000, GS1, CIS Controls Secure supply chains and ensure regulatory and ethical conformance.
Emerging AI-related ISO/IEC JTC 1/SC 42 Address risk, transparency, and trust in AI systems.
Standards Development and Extension Pathways. Standards in SE typically evolve through two complementary pathways. The first is the formal route, followed by standards development organizations (SDOs) such as IEEE and ISO/IEC, which emphasize stability, consensus, and global legitimacy. This approach is characterized by well-defined, gated stages including project authorization, working group (WG) formation, draft development, public review, and final balloting (International Organization for Standardization, 2023). These stages typically span 18–48 months for IEEE standards and 24–36 months for ISO/IEC standards (Tsui et al., 2022; IEEE Standards Association, 2023; International Organization for Standardization, 2023). The first version of IEEE POSIX standard, for example, progressed through this process via multiple iterations and ballots before becoming ISO/IEC 9945 (Isaak, 2005).
The second pathway is an upstream, community-led extension model, in which specifications evolve collaboratively in open WGs, consortia, or technical bodies before entering the formal process. This “implementation-first” or “parallel” model (Blind and Böhm, 2019) underpins the evolution of major standards: POSIX is maintained by the Austin Group prior to IEEE and ISO ratification (The Open Group, 2022); amendments to IEEE 802.11 (Wi-Fi) are incubated in dedicated task groups (IEEE 802.11 Working Group, 2023); and SPDX itself matured within the Linux Foundation community before being standardised as ISO/IEC 5962:2021 (International Organization for Standardization and International Electrotechnical Commission, 2021). Such upstream-first approaches offer advantages such as faster iteration, broader participation, and validation against real-world practice. These factors are critical in fast-evolving domains such as AI. Our work deliberately adopts this second model.
2.2. Software Bill of Materials (SBOMs)
From BOM to SBOM. The concept of a Bill of Materials (BOM) originated in manufacturing as a structured inventory of components and sub-assemblies within a product (Jiao et al., 2000). Applying this principle to software, a Software Bill of Materials (SBOM) provides a machine-readable inventory of software components, enabling organisations to understand what software they run, where it originates, and how it can be trusted (Xia et al., 2023b). SBOMs have become central to software supply chain security as modern systems increasingly depend on third-party and open-source components (Stalnaker et al., 2024).
SBOM standards. Three SBOM standards dominate the landscape: SPDX (Foundation, 2025), developed under the Linux Foundation and standardized as ISO/IEC 5962:2021 (International Organization for Standardization and International Electrotechnical Commission, 2021); CycloneDX (OWASP, 2025), introduced by OWASP and standardized as EMCA-424 (Ecma International, 2024); and SWID Tags (ISO/IEC, 2015), maintained by the U.S. National Institute of Standards and Technology (NIST) to provide transparent software identification. While SPDX and CycloneDX are both widely adopted and offer comparable core capabilities, their selection in practice is typically guided by organizational priorities. SPDX is often preferred for license compliance, provenance tracking and governance use cases, while CycloneDX is preferred for vulnerability management, security automation, and CI/CD integration (National Telecommunications and Information Administration (2021), NTIA; Cybersecurity and Infrastructure Security Agency (2023), CISA; OASIS Open, 2023).
Limitations for AI Systems. Despite their value, current SBOMs standards do not capture the full complexity of AI systems (Muirí, 2019). New artifact types, such as datasets and models, introduce deeper provenance, explainability, and compliance requirements. Dataset quality and lineage directly influence performance and fairness (Lu et al., 2022; Rajbahadur et al., 2021); model opacity necessitates metadata on interpretability, hyperparameters, and operational constraints (Barclay et al., 2023); and iterative lifecycles, including pre-training, fine-tuning, and redeployment, create intricate provenance chains and evolving compliance obligations (Hassan et al., 2024). These limitations have led both researchers and practitioners to call for extensions to SBOM standards that address the specific needs of AI systems, laying the foundation for the concept of an AIBOM.
3. Methodology
We followed a participatory and iterative approach to develop the AIBOM specification in the open, engaging multiple, diverse stakeholders. Figure 1 illustrates the process we followed.
Figure 1. Overview of AIBOM standard creation process and multi-faceted validation.
3.1. Working Group Formation
Motivation and scope. The AI Working Group (WG) was formed in 2021 under the SPDX community to explore how existing SBOM standards could be extended to represent AI systems. Its primary goal was to determine whether SBOMs captured the artifacts and metadata needed for regulatory compliance, license obligations, auditability, and transparency in AI systems.
Recruitment and participation. From the outset, the WG held weekly one-hour virtual meetings open to anyone interested. Early participation (2021 to mid-2022) relied on direct invitations and word-of-mouth recruitment through professional networks. To broaden engagement, the group launched a public mailing list on 7 June 2022, allowing contributors to sign up independently and join ongoing discussions.
Structure and operation. The WG ran open agendas, made decisions by consensus, and used mailing lists and GitHub for asynchronous discussion and review. Participants included scientists, researchers, professors, CTOs, product managers, AI and software developers, and legal and licensing experts.
Outcome 1 By May 2024, shortly after the official release of SPDX 3.0, the WG had held 82 meetings with 92 unique participants, 20 of whom attended more than six sessions. All meetings since April 2022 were recorded and archived publicly for transparency and traceability (contributors, 2024).
3.2. Defining the AIBOM Specification
Initial field definition. The WG’s first major task was to identify the core information an AIBOM specification should capture. Drawing on their diverse expertise and relevant research, participants proposed an initial set of fields during weekly meetings in early 2022. The objective was to ensure that the AIBOM could represent essential aspects of AI systems including datasets, models, and their associated metadata. At this stage, the focus was on brainstorming and cataloging all potentially relevant fields.
Incorporating existing practices. From July 2022, once a comprehensive list had been drafted, members systematically analyzed established documentation artifacts such as Model cards (Mitchell et al., 2019), Datasheets (Gebru et al., 2021), and AI factsheets (et al., 2019) to refine and expand the field set. The goal was not exhaustive mapping, but to identify the most relevant and widely applicable fields recognized as useful and appropriate in the AIBOM context. A detailed mapping of the fields from these sources that were incorporated, excluded and the reasons we did so, is provided in our whitepaper (Bennet et al., 2024).
Balancing adoption and completeness. Throughout this process, the WG prioritised adoption over comprehensiveness by defining a minimal set of required fields—those most likely to exist in real-world projects and satisfy regulatory or auditing needs. Each field had to meet strict inclusion criteria: (1) relevance to AIBOM goals, (2) availability of a representation method, and (3) consensus on its necessity. More ambitious metadata elements were intentionally deferred to future releases. In line with NTIA (National Telecommunications and Information Administration (2021), NTIA) recommendations and emerging best practices, the WG focused on a small, readily adoptable set of required fields. These fields were refined continuously until May 2024 through iterative review and discussion.
Outcome 2 The WG evaluated 103 candidate fields derived from existing tools and community proposals. The final specification defined 36 fields: 20 for the AI profile (five required) and 18 for the Dataset profile (six required).
3.3. External Feedback and Public Consultation
Presenting drafts and gathering early feedback. WG members frequently presented draft versions of the AI and Dataset profiles (that form the AIBOM specification along with the rest of the SPDX profiles, please see Section 4 for more details) at several formal venues, including Open Source Summit Europe (2022, 2023), Open Source Summit Japan (2022), and Open Source Summit North America (2023, 2024, 2025) and semi-formal WGs including the OpenChain Licensing WG (Coughlan, 2022), OpenSSF AI WG (Open Source Security Foundation (2025), OpenSSF), and IEEE P7014.1 WG (IEEE P7014.1 Working Group (2024), EEEPG). These events brought together open-source experts, software practitioners, and legal professionals, whose feedback was then discussed in WG meetings and incorporated into subsequent revisions of the specification.
Public release candidates and community review. Starting on 8 May 2023, the WG initiated a year-long public consultation process, releasing two versions of the AIBOM specification: two release candidates (RC1 and RC2) (spd, 2025). The SBOM tooling community was invited to review the specification, model, and profiles and to submit proposed changes as pull requests (PRs) to the public repository. In total, 20 PRs specific to AIBOM were submitted, 16 of which were accepted after detailed discussion in WG meetings.
Integration with SPDX. In parallel with the release candidate process, the WG collaborated with the broader SPDX project to ensure that the proposed extensions were incorporated and could work seamlessly with other elements of the SPDX specification. Since several new fields required changes to the underlying data model, the group regularly presented proposed modifications to the SPDX Technical Team (responsible for the specification’s architecture and publication). These discussions led to coordinated updates across all SPDX WGs, enabling seamless integration of the AIBOM related profiles and fields into SPDX 3.0 prior to its final release.
Outcome 3 After multiple rounds of refinement, SPDX 3.0, including the AI and Dataset profiles (please see Section 4 for more details) that encapsulate the AIBOM specification was officially released on 16 April 2024. The release was accompanied by a Linux Foundation blog post (Foundation, 2024) and a detailed whitepaper (Bennet et al., 2024) to support adoption and provide implementation guidance to the broader community.
4. AIBOM
Figure 2. Overview of SPDX profiles and their extensions for AI and Dataset.:
Building on the methodology described in Section 3, the working group integrated two new profiles:Dataset and AI, into the SPDX 3.0 specification. These extensions expand the SBOM standard beyond traditional software components to capture the artifacts underpinning AI systems. Figure 2 shows how the various SPDX profiles together represent an AIBOM specification (thereby extending the SBOM standard to represent AI systems. We refer to this as either AIBOM standard or AIBOM specification in our paper.
Designed to support regulatory compliance, license attribution, provenance tracking, and responsible development, they remain fully compatible with existing SPDX tooling and workflows. Detailed overview of our AIBOM specification is available in the SPDX 3.0 whitepaper (Bennet et al., 2024) and public repository. [^1] https://github.com/spdx/spdx-3-model As this paper focuses on the process of creating the AIBOM specification rather than the specification itself, we provide only a high-level summary below.
Profiles and extensibility in SPDX. SPDX 3.0 adopts a modular, profile-based architecture that enables domain-specific extensions while preserving interoperability. A profile defines a coherent subset of the model: the Core profile establishes foundational classes and vocabularies, while additional profiles extend these capabilities to represent licensing, build metadata, security, and software composition.
Within this framework, the AIBOM specification for an AI system is realized through the AI and Dataset profiles working alongside existing components (though they can also be used independently to document models or datasets). Rather than existing as standalone artifacts, these profiles extend SPDX in a modular way without fragmenting the broader ecosystem (Figure 2). They build on existing classes and relationships from the Core, Licensing, and Software profiles while introducing domain-specific metadata essential for describing AI artifacts. Together, they allow datasets and models to be represented as first-class supply-chain elements within the same SBOM framework.
AI profile. The AI profile captures AI model-specific components, capturing provenance (identifiers, versions, and licenses), architecture (type, domain, and evaluation metrics), training details (data sources, configurations, and resource usage), and model risks (limitations, safety considerations, and compliance evidence). It provides a comprehensive view of how models are developed, deployed, and maintained.
Dataset profile. The Dataset profile standardizes how datasets are described, capturing provenance (origins, sources, suppliers, and collection processes), descriptive details (size, modality, structure, and preprocessing), intended use (applications and purposes), and risk information (biases, sensitive data, and anonymisation methods). This enables traceability, accountability, and compliance throughout the AI lifecycle.
5. Validation
To demonstrate that our proposed AIBOM specification is comprehensive, practical, and fit for purpose, we designed a multi-faceted validation strategy. This process evaluated our specification from four distinct but complementary angles which we detail in the following subsections.
5.1. Regulatory and Standards Alignment
Table 2. Overview of key standards and regulations considered for AIBOM alignment
Studied Standard/Regulation Objective of the Standard/Regulation
EU Artificial Intelligence Act (EU AI Act) (European Parliament and Council of the European Union, 2024) Ensure safe deployment, fundamental rights protection, and risk-based regulation of AI.
EU Medical Device Regulation (EU MDR) (European Parliament and Council of the European Union, 2017) Demonstrate conformity and document risk management for medical devices placed on the EU market.
US Food and Drug Administration (FDA) Guidance (Center for Devices and Radiological Health, 2025) Support pre-market evaluation of software-based medical devices and promote cybersecurity risk management.
IEEE 7000 Series (Spiekermann, 2017) Promote ethical design, transparency, bias mitigation, and data governance in autonomous and intelligent systems.
Motivation. It is essential that our AIBOM specification captures all metadata required to verify compliance with key regulations and standards governing AI systems. Practical adoption depends on whether a supply chain standard enables Open Source Program Offices (OSPOs) and compliance auditors to fulfill their legal and governance obligations. In this section, we validate whether the fields defined in AIBOM specification can represent the information demanded by widely adopted and emerging standards. We focus on the EU AI Act, U.S. and EU medical device regulations, and the IEEE 7000 series, as they collectively span regulatory, safety-critical, and ethical requirements across diverse domains. Table 2 summarises the studied standards and regulations. Due to their extensive scope, we focused our analysis on the most relevant clauses and information elements that directly impact software supply chain compliance.
Methodology. We analyzed the alignment between fields captured in the AIBOM specification and the regulatory and ethical requirements defined by the standards and regulations shown in Table 2. The analysis was performed by the fourth and fifth author, who have extensive experience in standards development and regulatory alignment, and was further independently reviewed by first and the sixth authors for additional validation.
EU AI Act., we examined the obligations defined in Articles 49 and 60, which govern mandatory registration in the EU high-risk AI database prior to deployment or testing. Required information was grouped into four categories: Identification, System Details, Verification Details, and Application Details, comprising a total of 14 subcategories (e.g., provider contact, intended purpose, classification, testing plan, and involved parties).
US and EU medical device regulations, we analyzed our AIBOM specification’s alignment with Article 10 of Regulation (EU) 2017/745 and FDA cybersecurity guidance. Information elements were organized into three categories: package details, model details, and data details.
IEEE 7000 series, we evaluated over 40 subclauses from eight standards (7000, 7001, 7002, 7005, 7007, 7009, 7010, and 7014), covering key areas such as transparency, privacy, data governance, and algorithmic bias.
Results. Our analysis shows that the SPDX 3.0 AIBOM specification provides strong coverage of regulatory, safety, and ethical requirements. We found that we could successfully represent 13 of 14 information obligations under the EU AI Act, capturing all required elements from US and EU medical device regulations. Our validation shows that AIBOM specification aligns with over 40 subclauses across eight IEEE 7000 series standards i.e., more than 90% overall coverage indicating practical readiness for compliance adoption.
However, the analysis also revealed limitations that highlight opportunities for future refinement. While AIBOM captures most metadata required by the EU AI Act, it cannot yet explicitly represent the “parties involved in testing” relationship mandated by Annex IX (2), even though contact details can be documented in existing fields. Similarly, the strong alignment with medical device regulations is partly due to their reliance on free-text “summary” and “description” fields, which map to generic SPDX constructs such as informationAboutApplication and comment but may limit traceability and automation. Finally, although more than 40 IEEE 7000 subclauses on transparency, privacy, data governance, and algorithmic bias are covered, areas such as child and student data governance, robotic nudging, and environmental and social governance remain outside the scope of the current AIBOM specification. A detailed field-level mapping of these standards and their alignment with SPDX 3.0 is available in the official white paper (Bennet et al., 2024).
5.2. Validation against Key Industry Use Cases
Motivation. While regulatory alignment ensures that our AIBOM specification meets legal and compliance obligations, its practical value ultimately depends on how well it addresses real-world needs identified by industry stakeholders. In June 2025, the SBOM for AI Tiger Team (SBOM for AI Tiger Team, 2025), convened by the U.S. Cybersecurity and Infrastructure Security Agency (CISA) [^2] https://www.cisa.gov/, accessed September 30, 2025, articulated six foundational use cases that any AI-focused SBOM standard must support. These use cases span the full lifecycle of AI system deployment: regulatory compliance, Vulnerability and Incident Management, legal and intellectual property assurance, third-Party AI Risk Management, Open Source Model Risk Assessment and Model Lifecycle and Asset Management. A detailed overview of these usecases can be found here (Use, 2025).
Validating our AIBOM specification against these use cases is therefore essential to demonstrate that the specification is not only standards-compliant but also operationally relevant, usable by organizations in practice, and capable of supporting the core security, legal, and governance workflows envisioned by the broader AIBOM consumer ecosystem.
Methodology. The CISA “SBOM for AI” use cases are intentionally high-level; however, assessing whether our AIBOM specification is capable of capturing the metadata outlined in them requires concrete, testable requirements. We therefore operationalized these use cases by consolidating them into broader themes, assembling a focused evidence base from recent studies, extracting atomic requirements, and mapping them to SPDX 3.0 AIBOM fields before performing external validation.
Step 1: Consolidate use cases into themes (with rationale). Because several of the six use cases overlap conceptually and are instances of broader goals, we collapsed them into three themes to avoid double-counting and align with the structure of available evidence: (i) Compliance (Compliance; Legal & IP Protections), (ii) Open Source Risk Management (Vulnerability/Incident Management; Open Source Model Risk Assessment), and (iii) Supply Chain Management (Third-Party AI Risk Management; Model Lifecycle & Asset Management).
Step 2: Build an evidence base per theme. For each theme, we identified relevant research papers and technical reports that analyze concrete instances of the use cases. For example, Rajbahadur et al. (Rajbahadur et al., 2021), which outlines methods for conducting dataset license compliance analysis, was mapped to the Compliance category. All selected studies are listed in Table 3.
Step 3: Extract atomic requirements. The third author systematically reviewed each source and extracted concrete, verifiable requirements specific to its theme (template: requirement statement, evidence snippet/page, rationale). For example, Rajbahadur et al. (Rajbahadur et al., 2021) examine licensing and provenance risks in using public datasets for commercial AI; from this, we derived the requirement to capture original data sources and licensing/redistribution constraints for each dataset and derivative build. This process was repeated for all sources under each theme.
Step 4: Map requirements to AIBOM. We created a traceability matrix linking each extracted requirement to specific SPDX 3.0 AIBOM fields, indicating whether and how the requirement could be addressed by existing fields.
Step 5: Internal and external validation. The first author independently verified the extractions and mappings. We then presented the traceability matrix and representative examples to the AI SBOM Tiger Team (29 Sept 2025, attended by 12 participants) and the SPDX AI and Dataset WG (attended by four participants). Authors of this paper who were part of the working group did not participate in the validation exercise. Feedback was solicited on whether AIBOM could be effectively operationalized to address the six use cases. Meeting minutes from both working group sessions are publicly available. [^3] SBOM for AI (AIBOM) Tiger Team working group meeting minutes https://docs.google.com/document/d/1IpXG7XBOJnPl_hwFf3JZkDaFb0k2CnI0/edit?usp=sharing&ouid=110194678381965933391&rtpof=true&sd=true
Results. Our analysis demonstrated that the AIBOM specification could represent all 46 distinct requirements extracted from the eight studies reviewed across the three consolidated use-case categories. Table 3 summarizes the number of AIBOM fields across the AI, Dataset, and other SPDX profiles that satisfied the requirements identified in each study. In several cases, a single requirement could be addressed by multiple fields, underscoring the flexibility and composability of our AIBOM specification. A detailed traceability matrix, including the full set of extracted requirements and their field-level mappings, is available in our extended technical report (Rashno, 2025).
External validation. Feedback from participants in both the SPDX AI & Datasets and AI SBOM Tiger Team working groups was broadly positive. Reviewers agreed that the methodology and resulting mappings were sound and well-aligned with practical needs, while noting that additional time would be required for an exhaustive, line-by-line review. As our primary objective was to identify any major conceptual gaps or misinterpretations rather than obtain final endorsement, this feedback was encouraging. Both groups have committed to conducting deeper evaluations in future iterations, and we intend to incorporate any outcomes available before the camera-ready version of this paper.
Table 3. Alignment of consolidated AI SBOM use cases and the number of mapped fields across the AI, Dataset, and other SPDX profiles.
Categories References AI profile Dataset profile Other profiles
Compliance (Use, 2023a) 18 7 8
(Rajbahadur et al., 2021) 5 11 3
(Longpre et al., 2023) 4 3 3
Open Source Risk Management (Use, 2023b) 6 5 4
(Jiang et al., 2024) 3 8 4
(Stalnaker et al., 2025) 5 4 2
Supply Chain Management (Keskin et al., 2021) 7 3 3
(Schlegel and Sattler, 2023) 5 6 3
5.3. Validation based on Practitioner Interviews
Motivation. While the design of our AIBOM specification was driven by a large, multi-stakeholder working group, it was essential to validate the resulting fields through an independent lens. To ensure that our specification was both understandable and practical beyond the community that developed it, we conducted semi-structured interviews with practitioners who were not involved in the AIBOM working groups. Their perspectives served as an external “sanity check”, allowing us to assess whether the proposed fields align with real-world needs, uncover overlooked requirements, and identify potential barriers to adoption in industry practice.
Methodology. It is important to note that this validation was conducted prior to the official release of the AIBOM standard. This timing ensured that participants focused on the fields they considered essential for an effective AIBOM, rather than merely commenting on the limitations of our finalized specification. A detailed overview of our methodology is presented below.
Participant Recruitment. For this study, we recruited practitioners with experience working with SBOM or developing software with an AI component. We relied on our professional networks to recruit participants for this study. Throughout the 2022-mid 2023 period, we directly emailed experts who are known to have experience with SBOMs or developing AI software. From the individuals with relevant experience who responded to our emails, we ultimately selected ten participants for our study.
Table 4 shows the demographic details of the ten participants. The interview participants varied in their work experience, work domain, and their future relationship with AIBOM.
Part. ID Software/AI Dev./SBOM Experience (yrs) Domain AIBOM Role
P1a 11 / 9 / 0 Entertainment Software Consumer
P2b 25 / 0 / 4 Healthcare Software Consumer
P3c 4 / 2 / 1 ICT Consumer
P4c 14 / 2 / 2 ICT Both
P5d 15 / 6 / 0.5 Academia Consumer
P6e 20 / 0 / 16 Open Source Foundation Both
P7f 4 / 3 / 0 Logistics Software Creator
P8g 12 / 12 / 0 IT Consultancy Both
P9d 9 / 5 / 0 Software Research Both
P10h 15 / 0 / 5 IT and Networking Both
  • a
    Lead Data Scientist
  • b
    Chief Operating Officer
  • c
    SE Researcher
  • d
    Research Scientist
  • e
    General Manager
  • f
    ML Engineer
  • g
    Consultant
  • h
    SBOM Advocate
Table 4. Study Participant Demographics and Experience
Interview Protocol. First, we drafted an interview questionnaire and presented it to two experts from academia and industry. Based on their feedback, we finalized the interview questions. The 18 questions in the interviews centered around the following themes: the participants’ past experience with AI and software development, a working definition of AI software/system traceability, current practices and tools available for ensuring AI system traceability, and expectations from a new AIBOM specification.
We conducted semi-structured interviews with each participant as it allowed for improvisation and exploration of the studied objects (Wohlin et al., 2012). All the interviews were conducted via video calls and lasted 30 to 60 minutes each. To avoid bias, we did not show the proposed fields to the participants. We received consent from all participants to record the interviews and use the recordings for research purposes. We transcribed the recordings using software and manually checked for errors.
Transcript Coding and Analysis. First, two authors sat together to read through two interview transcripts in full and induce codes from interview passages. Then, the authors independently read the remaining interview transcripts thoroughly and assigned codes to passages. Next, the authors met again to filter out the codes irrelevant to our goal and grouped the independently identified codes for each transcript. Once a consensus was reached about the codes and categories from the interviews, we computed the overlap between the fields included in our AIBOM specification and the concepts proposed by the interview participants.
Table 5. AIBOM fields validated by interviews.
Category Proposed Field SPDX 3.0 AI/Dataset Fields
Data Lineage originatedBy, dataCollectionProcess, datasetUpdateMechanism, downloadLocation
Provenance downloadLocation, originatedBy
Metadata spdxId, name, packageVersion, buildTime, releaseTime, primaryPurpose, datasetSize
Assumptions datasetNoise
Preprocessing dataPreprocessing, anonymizationMethodUsed
Licenses hasConcludedLicense, hasDeclaredLicense
Intended use intendedUse
Personal info hasSensitivePersonalInformation
Models Behaviour metric, metricDecisionThreshold, safetyRiskAssessment, modelExplainability
Parameters hyperparameter
Algorithms informationAboutTraining, typeOfModel
Intended use primaryPurpose, informationAboutApplication
Assumptions informationAboutApplication, limitation
Biases knownBias, limitation
Licenses hasConcludedLicense, hasDeclaredLicense
Features modelExplainability, domain, typeOfModel
Code License hasConcludedLicense, hasDeclaredLicense
Environment Hardware sensor, energyConsumption, inferenceEnergyConsumption, trainingEnergyConsumption
Configuration proposed in SPDX 3.1
Deployment proposed in SPDX 3.1
Process Data Collection dataCollectionProcess, sensor
Training informationAboutTraining, trainingEnergyConsumption, finetuningEnergyConsumption
Governance Dependencies contains, downloadLocation
Attribution originatedBy, suppliedBy, name, spdxId
Ethical considerations safetyRiskAssessment, knownBias, useSensitivePersonalInformation, hasSensitivePersonalInformation, confidentialityLevel, anonymizationMethodUsed, standardCompliance, intendedUse
Design decisions & limitations limitation, modelDataPreprocessing, dataPreprocessing, metricDecisionThreshold, typeOfModel, hyperparameter
Results. With the exception of Configuration and Deployment-related fields (which we had deliberately earmarked for inclusion in a future version due to difficulty with capturing them) all fields proposed by participants could be represented within the current SPDX AIBOM 3.0 specification. Table 5 presents the mapping between the fields suggested by study participants and those included in the AIBOM specification.
The primary theme emerging from the interviews was the widespread inadequacy of current traceability practices for AI systems, which participants described as ad hoc and insufficient for ensuring proper governance. This sentiment was powerfully articulated by an executive director (P2), who stated, ”I think it’s a general mystery across the industry… I don’t think that [traceability] adequately exists within the AI community around algorithms and training sets to be able to demonstrate that traceability from an auditing standpoint.” This lack of a standardized, auditable record is exacerbated by real-world development pressures. A machine learning engineer (P7) from a startup described this vividly: ”It’s something that we lack and it’s because… the general nature of a startup, we usually deprioritize documentation and it usually also ends with a lot of pain for us… because we don’t have the documentation, I usually spend so much time trying to explain anyway.”
Beyond identifying these gaps, participants also provided clear insights into what an effective traceability solution should capture. When asked what information they need when consuming a pre-trained model, a lead data scientist (P1) offered a detailed wishlist: ”What do I look for? I look for license… support… When, which training data was used, what demographic was used, and what biases do they have? …And these are stuff that you want captured in the document, because that’s what I’m looking for [as a model consumer]… what’s the [reported] accuracy? How did you test it?” This practitioner’s desired list of artifacts, spanning licensing, data provenance, bias, and evaluation metrics, directly aligns with the core categories of the proposed AIBOM framework, confirming that its design addresses the expressed needs of practitioners. These insights will continue to guide the future evolution of our AIBOM specification.
5.4. Field Validation in Industrial Setting
We report an external study conducted at a large multinational software organization to evaluate AIBOM’s suitability for representing AI system documentation use cases in practice. The study was conducted by a team of security architects, data scientists, DevOps experts, and AI developers. From June 2024 to September 2024, the team members worked on identifying the overlap between the fields available in our AIBOM specification and the fields encountered by practitioners during the organization’s development of an AI system.
Methodology. The organization already had a checklist for developers to evaluate how their AI software development affects the internal ethics policy. Moreover, the legal department had a checklist to assess the legal impact of each AI-related software release. The researchers investigated how many of the fields in these checklists could be populated if the organizations already had AIBOMs constructed as part of the AI software development process. Furthermore, they investigated whether our AIBOM specification can be used to automatically generate model cards. Then, they investigated to what extent AIBOM generation for third-party models can be automated using web scraping or external APIs, such as Hugging Face (Hugging Face, 2025).
Results. The analysis found that all of the fields required in internal checklists in the organization’s AI system development process can be populated from the AIBOM specification’s fields, demonstrating the comprehensiveness of AIBOM in practice. They observed that 60% of the fields in the existing model cards could be extracted using the fields in the AIBOM specification. Furthermore, they found that external APIs and web scraping could populate 40% of the fields when generating AIBOMs for third-party models used at the organization. This demonstrated that AIBOM generation could be partially automated for practical use cases, reducing the manual effort needed.
6. Standards Development as Action Research
A key contribution of our paper is demonstrating how the development of a complex, multi-stakeholder standard can be systematically understood through the lens of Action Research (AR). AR is a methodology that addresses real-world problems while producing actionable knowledge (Baskerville, 1999; Elden and Chisholm, 1993; Ferrario et al., 2014). By retrospectively framing our large-scale effort to extend SPDX and create an AIBOM specification as an AR process, we provide a concrete example of how community-driven standards development can follow a disciplined, research-grounded approach in rapidly evolving domains.
Although we did not initially conceive this project as AR, its conduct and outcomes closely align with AR principles. AR is fundamentally concerned with changing practice, with an “explicit aim to act in the real world and change the state of practice.” Our goal was to create a functional, community-driven standard for AI transparency and compliance i.e., an intervention by design. We therefore frame our work using the canonical AR cycle (Staron, 2020) to structure and reflect on the process.
Our AR framing is particularly relevant in software engineering (SE), where controlled experiments, case studies, and surveys dominate, while AR remains comparatively rare (Petersen et al., 2014). Sjøberg et al. (Sjøberg et al., 2007) observe that part of the reason is a limited understanding of AR’s role in SE. By analyzing the AIBOM specification development effort through AR lens, we offer a concrete example of its application in the wild to the software engineering body of knowledge.
Table 6 maps the main steps of our methodology to the canonical AR stages. This perspective highlights the key enablers of our success: continuous stakeholder engagement, iterative adaptation, and feedback-driven evolution. Further, it demonstrates how standards development can operate simultaneously as a scientific method and a practical intervention.
Table 6. Mapping AIBOM Development to the AR Cycle
AR Phase AIBOM Development Stage
Diagnosing Identifying a practical problem WG formation and scoping: Addressed the lack of a standardized way to describe AI components for traceability, compliance, and transparency.
Action Planning Designing an intervention Field definition and framework design: Extended SPDX to include AI artifacts by analyzing model cards, datasheets, and existing practices.
Action Taking Implementing the intervention Drafting and presenting profiles: Created draft AI and Dataset profiles and presented them at major forums to gather feedback.
Observation Evaluating impact Feedback collection: Collected input via release candidates, pull requests, and discussions to assess utility and design.
Reflection Learning and refining Profile refinement: Iteratively updated profiles based on feedback, aligning them with real-world adoption needs.
Iterative Cycles Repeating and improving Release iterations: Integrated feedback into successive release candidates, refining the specification before final release.
Knowledge Dissemination Sharing outcomes Publishing and outreach: Released SPDX 3.0 with AI and Dataset profiles, supported by a whitepaper, blog posts, and documentation.
7. Lessons Learned and the Road Ahead
This section first distills the key lessons learned from our experience developing the AIBOM standard in an open, multi-stakeholder setting, and then outlines the road ahead by discussing how these lessons inform future extensions, particularly in the context of foundation model-powered software.
7.1. Lessons Learned
Present Work-in-Progress Early and Solicit Feedback Often. Public presentations of early drafts, in formal venues such as OSS Summit sessions (see Section 3.3) and other WGs provided critical feedback between mid-2022 and 2024. This early engagement not only strengthened the evolving standard but also built practitioner trust and buy-in. Feedback from outside our WG often prompted significant design changes. For instance, input from the OpenChain WG led to the introduction of a standardsCompliance field to capture which standards an AI model or dataset already adheres to, which is a key requirement for audits. A question during an OSS NA Vancouver panel about the queryability of fields such as energyConsumption and metrics prompted us to replace free-form text with structured types. Similarly, discussions at OSS Summit Japan 2022 revealed that tracking AI and data lineage separately would complicate adoption, leading us to merge them into a unified dataCollectionProcess field. Based on our experience, in line with previous studies (Centre., 2019), we suggest that continuous, feedback-driven development is not merely advantageous but essential for open standards in fast-moving domains like AI.
Prioritize Adoption Over Comprehensiveness.  Early drafts that attempted to capture every conceivable detail of an AI system consistently faced pushback from practitioners. Most organisations simply do not maintain information at that level of granularity, and a standard that demands it becomes impractical (Jespersen, 1995; Isaak, 2005). A similar pattern is evident with model cards, which prescribe numerous fields, and as a consequence, over 60% of models and 70% of datasets on Hugging Face lack a complete model card (Stalnaker et al., 2025; Oreamuno et al., 2024). Several prior studies document similar resistance to heavyweight standards (Pfleeger et al., 2002; Schmidt, 2000).
We therefore optimized our AIBOM specification for adoption by defining a small set of readily recordable required fields and enforcing strict entry criteria. In some cases, we intentionally excluded ambitious goals to improve practicality. For example, rather than attempting to enumerate every potential form of bias, we introduced a single knownBias field to capture only documented biases. Additional fields were deferred to future releases, enabling incremental evolution as practices mature.
We also found that two factors: metadata scale and availability, strongly influence adoption. For instance, a comprehensive AIBOM for a self-driving car could span several gigabytes, making generation, storage, and maintenance impractical. At the same time, much of the desired metadata simply does not exist in accessible form. Even foundational datasets such as ImageNet and CIFAR-10 do not fully disclose their data sources (Rajbahadur et al., 2021). Supporting fields that can be automatically collected or derived are therefore essential. Recognizing this, we are now collaborating with the SPDX Implementers WG (SPDX Project / Linux Foundation, 2025) and CISA SBOM-o-RAMA (Cybersecurity and Infrastructure Security Agency (2025), CISA) to improve automation support in future iterations. These experiences highlight a broader reality: premature attempts at exhaustive coverage can undermine adoption. While similar lessons have been echoed several decades ago during POSIX standardization (Jespersen, 1995; Isaak, 2005), they ring true even today. Prioritizing a minimal, usable core builds early momentum and a foundation for richer requirements as the ecosystem matures.
Plan for Churn and Conflict from Day One. In long-running, community-driven standardization efforts, both contributor turnover and conflicting priorities are inevitable, and planning for them early is essential. Several participants deeply involved in shaping the initial drafts disengaged as their organizational priorities shifted, while others joined later and made key contributions during specific phases. This irregular participation often left critical work unfinished, slowing progress. In response, we established a rotating core group with decision-making authority to ensure continuity while still enabling broader community contributions.
Both subtle and overt disagreements were also a constant reality. In a multi-stakeholder setting with competing interests, participants frequently proposed adding, removing, or rewording fields, often with vague or narrowly scoped motivations. To prevent these debates from stalling progress, we consistently applied the evidence-based acceptance criteria defined in Section 7.1. This process significantly reduced friction and kept the WG focused on shared priorities rather than endless debate. As the AI ecosystem continues to diversify, sustaining continuity amid organizational churn and aligning competing priorities will become even more challenging. Embedding these mechanisms early helps ensure that open standards remain both stable and adaptable over time.
7.2. Road Ahead
Evolving the AIBOM to Capture FM-Powered Software. The rapid mainstream adoption of foundation models (FMs) and FM-powered software (FMware) since mid-2023 has fundamentally changed what an AIBOM specification must capture. FMware represents a clear departure from earlier paradigms. Whereas traditional AI-powered systems, using simple machine learning models such as Support Vector Machines or Neural Networks, are largely static artifacts with well-defined data sources and predictable lifecycle boundaries; FMware is continuous, adaptive, and open-ended (Hassan et al., 2024). It introduces new asset types such as prompts and downstream fine-tuned variants. Agents, which orchestrate workflows and mediate decision-making across multi-model pipelines, become first-class entities whose roles and interactions require explicit documentation (Hassan et al., 2024; Rajbahadur et al., 2024). FMware architectures also add operational layers such as grounding (linking decisions to verifiable sources) and guarding (enforcing safety, policy, and compliance), each producing additional artifacts and dependencies that must be tracked (Rajbahadur et al., 2024). These shifts, coupled with increasing regulatory expectations and enterprise risk-management requirements, make the evolution of the AIBOM to capture FMwareBOM both necessary and urgent. This next iteration expands beyond static system representation to capture agent behaviors, orchestration context, grounding sources, guardrail policies, data lineage, and lifecycle transformations. While our current AIBOM can adequately capture traditional AI-powered software, as evidenced by our validation in Section 5, we need to evolve it urgently for FMware.
To address FMware’s expanded requirements, we are preparing an update to AI and Dataset profiles for an SPDX 3.1 release candidate (SPDX Working Group, 2025a) that introduces fields for agent identity and capabilities (SPDX Working Group, 2025c), richer provenance links connecting datasets to generation prompts and retrieval contexts (SPDX Working Group, 2025d), and mechanisms to capture adaptation events, fine-tuning lineage, and downstream modifications (SPDX Working Group, 2025b). These changes aim to ensure that FMwareBOM captures the full lifecycle of modern AI systems while remaining automatable, auditable, and compatible with evolving compliance practices.
Making the Future AIBOM Tool-Native and Automatable. As highlighted in our Section 7.1, tooling is essential for adoption, and as discussed above, FMware significantly increases the scale and complexity of what must be tracked. To address this, we are designing the evolved version of AIBOM to be tooling-native from the outset. Our goal is to ensure that key fields can be extracted directly from development pipelines or inferred through lightweight analysis, while still meeting regulatory and risk-management requirements. Our hope is that this approach will enable greater automation throughout the FMware lifecycle—from generating BOMs and validating schema conformance during CI stages to assessing compliance and policy obligations automatically. Together, these design choices aim to make AIBOM not only richer in representation but also a practical, automatable, and auditable part of the FMware development process.
Enabling Interoperability Across the AI Supply-Chain Ecosystem. A key focus of our roadmap is ensuring that the next generation of SPDX AIBOM does not evolve in isolation but becomes a connective layer across a fragmented ecosystem of complementary standards and initiatives. Currently, efforts such as SLSA (for provenance and build integrity) (OpenSSF / SLSA Project, 2025), OpenChain (for organisational compliance) (OpenChain Project / The Linux Foundation, 2025), Croissant (et al., 2024) and emerging security frameworks like Coalition for Secure AI (Coalition for Secure AI, 2024) all address critical but distinct layers of the AI supply chain, yet lack a unified integration point. To address this, we are actively working to align and interoperate with these complementary initiatives. For example, we are exploring how FMwareBOM can serve as the canonical machine-readable artifact for AI compliance within the OpenChain AI WG, collaborating with OpenSSF and CISA to integrate provenance, verification, and security metadata, and engaging directly with industry and open-source leaders through collaborative initiatives. Towards this goal, we organized a panel at Open Source Summit North America 2025 on harmonizing SLSA provenance and SPDX SBOM (Maruseac et al., 2025), soliciting community input on how the two standards can complement and strengthen each other. Our goal is to reduce duplication, enable seamless metadata flow across layers, and build a cohesive ecosystem where complementary standards reinforce one another. The SPDX 3.0 specification is being prepared for formal submission to ISO as part of the effort to update the ISO/IEC 5962 standard. This planned ISO standardization, which will involve reviews by member bodies and Technical Committees, is anticipated to significantly boost adoption and ecosystem growth.
8. Related Work
To contextualise our contribution, we present prior work on emerging approaches for trustworthy AI, the challenges of multi-stakeholder collaboration in software engineering, and the application of AR as an intervention methodology.
Emerging approaches for trustworthy AI. Multiple initiatives aim to improve AI system trustworthiness by enhancing auditability, traceability, compliance, and provenance. DataBOM treats datasets as first-class artifacts to enable provenance tracking (Liu et al., 2024), but provides limited model-level detail. CycloneDX ML-BOM extends SBOM practices to ML models and datasets (bac, 2024), yet overlooks lifecycle complexities. Structured documentation approaches like IBM AI Factsheets (bac, 2025), Model Cards (bac, 2022), and Datasheets (Gebru et al., 2021; Mitchell et al., 2019) improve reporting but lack standard compliance and holistic AI system level coverage. AI Cards (Golpayegani et al., 2024) capture richer risk information but require tooling support, while TAIBOM (NquiringMinds LTD, 2025) automates BOM creation with novel, ecosystem-incompatible schemas. Broader governance efforts, including the NIST AI RMF (AI, 2023), Microsoft’s Responsible AI Standard (Microsoft, 2022), and Croissant (et al., 2024), promote accountability but remain disconnected from software supply chain realities, with Croissant focusing primarily on dataset metadata.
Despite their contributions, these solutions remain fragmented, either treating artifacts like datasets or models in isolation or proposing high-level frameworks detached from implementation. This piecemeal landscape leaves practitioners without a single, machine-readable artifact offering an end-to-end view of an AI system’s composition and provenance. Our goal with AIBOM is to address this gap: by extending the dependency-aware structure of SBOMs, it unifies data, models, and lifecycle processes into a cohesive, auditable standard.
SBOM research in SE. SBOMs are widely used for vulnerability management, transparency, risk assessment, and supply chain integrity (Qvarfordt, 2024; Martini, 2023). Empirical work shows that they reduce remediation times by enabling rapid dependency analysis (Benedetti et al., 2024) and improve auditability through provenance and licensing metadata (Martini, 2023). However, their impact depends heavily on tooling completeness and metadata quality (Stalnaker et al., 2024; Xia et al., 2023a). Researchers and practitioners alike have proposed extensions, such as blockchain-enabled SBOMs for tamper resistance (Xia et al., 2024) and DataBOMs for capturing provenance in data pipelines (Barclay et al., 2019; Liu et al., 2024). However, our study is the first document the process of creating AIBOM specification in a global, multi-stake holder setting.
Standardization Experience Reports. Existing experience reports on standardization in SE primarily offer historical narratives rather than structured, reproducible approaches to extension. Classic retrospectives on POSIX trace its evolution from user-group initiatives to IEEE and ISO ratification, providing valuable lessons but limited methodological insight into how such extensions were conceived and executed (Isaak, 2005; Jespersen, 1995). Complementary procedural documents, such as the Austin Group’s SD/6, detail governance aspects like defect handling and new work item procedures but remain detached from practitioner perspectives (Josey, 2012). Similar gaps exist in other domains: for example, Accellera’s SystemVerilog report documents a successful upstream-to-IEEE transition yet does not codify the development methodology (Sutherland, 2002). Our work complements and extends this literature by providing a detailed, transparent account of the end-to-end experience of developing an AI-focused extension, capturing decision-making, iteration, validation, and stakeholder engagement and maps standards development as an AR cycle that can serve as a replicable blueprint for future standardization efforts.
Multi-stakeholder studies in SE. Software engineering is fundamentally a socio-technical endeavor, shaped by the interactions of diverse actors including developers, managers, customers, and regulators. Multi-stakeholder studies investigate this complexity, analyzing the conflicting goals and socio-technical dependencies that influence project outcomes. Prior work highlights the benefits of incorporating multiple perspectives for improving requirements elicitation (Davis, 1993) and risk management (Alqahtani et al., 2017), particularly in large-scale projects (Hoda et al., 2017).
However, this literature also documents significant challenges, such as balancing heterogeneous interests and coordinating communication across organizational boundaries (Ralph, 2016). These challenges are magnified in the context of open, community-driven efforts, such as standards development, which represent a distinct and highly complex form of multi-stakeholder collaboration. Unlike projects within or between firms, creating a standard involves a decentralized group of volunteers with no single organizational authority, making consensus-building and governance particularly challenging. Therefore, the goal of this paper was to document the high-complexity domain of multi-stakeholder collaboration using open standard development as a unique case study.
Action Research in SE. Given the complex realities of software development, AR has emerged as a key methodology for structuring industry–academia collaboration, enabling researchers to intervene in real-world settings and iteratively refine tools and processes (Avison et al., 1999; Petersen et al., 2014). Its maturity in SE is reflected in recent efforts to formalize methodological rigor (Staron, 2025) and hybridize approaches, such as Design Science Action Research (DSAR), which combine artifact construction with organizational change (de Castro et al., 2025).
However, prior work shows that AR is typically confined to single organizations or bilateral partnerships, focusing on internal process improvement (Dittrich et al., 2024). Its application to open, community-governed initiatives producing public standards remains largely unexplored. To our knowledge, this work is the first to apply AR in such a setting, adapting it from a tool for organizational change to a framework for coordinating a global, multi-stakeholder community around a shared standardization goal.
9. Conclusion
Our experience demonstrates that Action Research can serve as a powerful framework for guiding standardization efforts in fast-moving, multi-stakeholder domains. By structuring the development of AIBOM specification around iterative diagnose-design-evaluate-reflect cycles, we showed how open collaboration, continuous feedback, and evidence-based iteration can transform fragmented community efforts into a coherent, widely applicable standard. The lessons distilled from this process provide a blueprint for future initiatives like our evolution of AIBOM specification in SPDX 3.1. Looking ahead, we envision AR-driven approaches playing a central role in shaping agile, trustworthy standards that evolve alongside the technologies they aim to govern.
Acknowledgements. The participation of Arthit Suriyawongkul in the WG is made possible through the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6224 (Research Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real)). We express our sincere gratitude to Helen Oakley and Rhea Michael Anthony for detailing the validation they conducted with our AIBOM framework in their industrial setting, as reported in Section 5.4. Furthermore, we would like to thanks Prof. Daniel M. German and Prof. Zhen Ming (Jack) Jiang for their valuable feedback and mentoring. Finally, we would also like to acknowledge the valuable contribution of all the members who participated in our SPDX AI and Dataset Working Group meetings.
Disclaimer. Generative AI tools were used for copy-editing and table formatting. All experiments, analysis, and writing were performed by the authors, who thoroughly reviewed the content. This complies with IEEE and ACM policies on AI use in publications.